# A Test-Time Learning Approach to Reparameterize the Geophysical Inverse Problem with a Convolutional Neural Network

### v1

@@ Abstract
Regularization is critical in solving the ill-posed geophysical inversion problems. Explicit regularization is often used, but there are opportunities to explore the implicit regularization effect inherently from a Neural Network structure. Researchers in Computer Vision (CV) have discovered that the Convolutional Neural Network (CNN) architecture inherently enforces a regularization that is advantageous for addressing diverse CV inverse problems, including de-noising and in-painting. In this study, we examine the applicability of this implicit regularization to geophysical inversions. The CNN maps an arbitrary vector to the model space (e.g., log-conductivity on the simulation mesh). The predicted subsurface model is then fed into a forward numerical simulation process to generate corresponding predicted measurements. Subsequently, the objective function value is computed by comparing these predicted measurements with the observed field measurements. The backpropagation algorithm is employed to update the trainable parameters of the CNN during the inversion. Note that the CNN in our proposed method does not require training before the inversion, rather, the CNN weights are estimated in the inversion algorithm, hence this is a test-time learning (TTL) approach. The results demonstrate that the implicit regularization provided by the CNN can be useful in DC resistivity inversions. We also provide a detailed discussion of the potential sources of this implicit regularization and some practical guides for applying the proposed method to other geophysical scenarios. The proposed approach for reparameterizing the inverse problem can be adapted to other Tikhonov-style geophysical inversions.
@@
Convolutional neural network (CNN), deep image prior (DIP), deep learning (DL), direct-current resistivity inversion, and parameterized inversion (PI).
@@ I Introduction
Geophysical inversions are widely used in mining, environmental, and engineering applications [22], [25]. Because geophysical inversions are ill-posed, regularization is important for obtaining a geologically reasonable solution. A conventional approach employs the Tikhonov-regularized method where an explicit regularization term, comprised of smallness and smoothness, is included in the objective function [22]. The L2 norm is standard in the Tikhonov-regularized methods [3], but there is research on using different norms to promote sparse or compact structures [18]. Researchers also developed strategies for incorporating physical properties [28], [29] into the Tikhonov-regularized methods. The recent emergence of deep learning has garnered significant attention from researchers, particularly regarding the integration of machine learning algorithms into the regularization [14], [24], [25], [30].
@@
One main approach is to search for a solution in a latent space whose distribution contains prior information learned from the training samples [1]. For example, in the Ground-penetrating radar (GPR) inversion, Laloy et al. [6] and Lopez-Alvis et al. [9] trained Variational Autoencoders to map the distributions of the spatial patterns of the subsurface materials into simple distributions and integrated this learned prior into the objective function. In the seismic full-waveform inversion (FWI), Mosser et al. [10] trained a generative adversarial network with the knowledge of subsurface geology. These methods predict more realistic subsurface models. In the electromagnetic (EM) inversion problems, B. Liu et al. [7] proposed the method PhResNet [7], which reconstructs the subsurface model by a U-net. B. Liu et al. [7] stated that reconstructing the resistivity models with CNN is beneficial by showing the inversion results with a training dataset. Examining the potential benefits of reparameterizing the subsurface model without a training dataset is one motivating factor for our work. Both the utilization of generative models and PhResNet require a training dataset, which implies that we need prior knowledge of what we are looking for in the subsurface (e.g. the possible geological models or subsurface materials of the survey domain). Therefore, this approach utilizes a physically motivated prior.
@@
Another approach employs so-called test-time learning (TTL) methods which do not require any training data when applied to each case. Rather than adjusting the weights of a neural network using training samples, the weights of a neural network are learned at the time the neural network is applied to the inversion process. TTL was first applied in the inverse problems in computer vision (CV). One example of the inverse problem in CV is in-painting (e.g. filling in corrupted portions of an image). This is achieved by learning the self-similarity between the corrupted image and the predicted image which is the output of the neural network. Researchers have found that some neural networks, such as Convolutional Neural Networks (CNNs) and Graph Convolutional Networks (GCNs), inherently enforce a regularization that is advantageous for addressing diverse CV inverse problems or mesh restoration problems [4], [17]. TTL methods have been applied to many biomedical imaging problems later [13], [26]. In geophysics, TTL has been employed in seismic FWI inversion by integrating the ML architecture into the inverse process in different ways [5], [11], [12], [15]. He et al. [5] and Wu et al. [11] reparameterized the velocity model into the CNN domain and attributed the regularization effect to the prior information learning by fitting the initial velocity models. In contrast, Zhu et al. [15] stated that this regularization effect comes from the CNN architectures as what Ulyanov et al. [17] proposed. Zhu et al. [15] examined this statement in the synthetic cases and found that this regularization effect is beneficial in cases where the noise is high. Utilizing trainable weights within the CNN to parameterize the subsurface model worked well to regularize the seismic inverse problems, but the prior they impose on the solution is not a physically motivated prior.
The physics that governs seismic FWI is the wave equation, and the strategies taken to solve the wave equation differ from those used in EM and potential field inversions, which are the so-called Tikhonov-style inversions. In this study, we would like to know the difference of the EM inversion results between the conventional Tikhonov-regularized methods and the methods with implicit regularization. In this work, we choose the direct current (DC) resistivity inversion, which is one type of EM inversion, but the proposed method can be applied to other Tikhonov-style geophysical inversions. Moreover, we explore the source of this implicit regularization effect and state that it is partly from the bi-linear upsampling operation, which smooths the final predicted model efficiently in some scenarios. Throughout, we will refer to this method as the Deep Image Prior Inversion (DIP-Inv), as it was inspired by the work of [17], which introduced the concept of the Deep Image Prior in computer vision.
@@
We will begin with an overview of the methodology and describe the modifications we make to the pipeline of inversion. Then we will show the results using the DIP-Inv. The results illustrate that the CNN reparameterization provides useful regularization effects; therefore, DIP-Inv gives better inversion results than the conventional methods in some cases. After that, we explore different components of DIP-Inv and identify the bi-linear upsampling operation as one of the main sources of this implicit regularization effect.
@@
    Compare inversion results obtained from conventional Tikhonov-regularized methods with sparse norms and obtained from the DIP-Inv approach on multiple synthetic examples to illustrate that the implicit regularization effect inherently from CNN is useful in the EM inversions.
@@
    Examine some key choices in the CNN architecture to illustrate that the bi-linear upsampling operation is one of the main sources of this implicit regularization effect and the Dropout technique is also useful.
@@
    Discuss some practical details for connecting the forward simulation, written in SimPEG [2], with PyTorch [21].
@@ II Methodology
In this section, we will first compare the conventional method for solving the geophysical inverse problem and the proposed method (DIP-Inv). Then, we will illustrate the structure of DIP-Inv. Finally, we will explain the objective function of the proposed method in detail. Note that we only conduct 2D DC resistivity inversions in this study, but this methodology can be adapted to 3D DC resistivity inversions or other Tikhonov-style geophysical inversions.
@@ II-A Overview of the Conventional Inversion Method
$\boldsymbol{\nabla}\cdot\sigma\boldsymbol{\nabla}\phi=\boldsymbol{\nabla}\cdot\mathbf{j}_{source}$ (1)
where $\sigma$ is the electrical conductivity, $\phi$ is the electric potential, and $\mathbf{j}_{source}$ is the source current density. We solve this equation using a finite volume approach where we discretize the subsurface using a tensor mesh (c.f. [16]), which defines the 2D mesh space $M$. The mesh consists of a core region, with uniform cells, and padding regions on the lateral boundaries and the bottom of the mesh, with expanding cells to ensure boundary conditions are satisfied. We use SimPEG, an open-source package, to implement the forward modelling [2].
@@
The trade-off parameter $\beta$ weights the relative importance of the data-misfit term and the regularization term. Often, a $\beta$-cooling strategy is adopted, and the value of the trade-off parameter $\beta$ is gradually reduced as the inversion progresses [22]. The regularization term $\phi_{m}$ generally includes a smallness term which penalizes the difference from a reference model $m_{ref}$, and a first-order smoothness term that penalizes spatial gradients in the model. The relative influence of each of these terms is weighted by the scalar parameters $\alpha_{s,x,z}$. The standard choice of norms in the regularization is $p_{s,x,z}=2$, however, different norms where $p_{s,x,z}\leq 2$ can be chosen, and we will use the implementation described in [18] for these inversions. Optionally, sensitivity weights can be included in the weight matrices $W_{s,x,z}$, and these have been shown to help improve the convergence of inversions where $p_{s,x,z}\leq 2$, [18].
@@
II-B Overview of the Proposed Parameterized Inversion method-DIP-Inv
Figure 1: Proposed parameterized inversion method. Compared with the conventional method, DIP-Inv reparameterizes the mesh space into the CNN-weighs space.
We are interested in exploring if implicit regularization can be advantageous in solving the inverse problem, so the main update that we make to the inverse problem posed in 2 is that we parameterize the model in terms of a CNN-based function.
In our proposed method, a CNN-based function $L$, whose weights are denoted as $w$, maps an arbitrary vector $z\in R^{k}$ to a model $m$ in the mesh space $M\in R^{W\times H}$ (See Fig. 1).
@@ II-B Overview of the Proposed Parameterized Inversion method-DIP-Inv
Note that the padding cells are included in $m$. The predicted subsurface model $m$ is then fed into the forward numerical simulation to generate the corresponding predicted measurements $d^{pred}$. The modified optimization problem that we aim to solve is
@@
Compared with the supervised end-to-end machine learning methods in the EM inversion [27], DIP-Inv imposes physical constraints in the forward simulation. Compared with the conventional Tikhonov-regularized methods, the objective function in the proposed method doesn’t have the smoothness term since the smoothness constraint is implicitly included in the upsampling operator in $L$ (section IV-A). Also, we remove the sensitivity weighting from the regularization without deteriorating the inversion results. The backpropagation algorithm, which is implemented by PyTorch, is employed to update the trainable parameters $w$ in function $L$. Since the weights $w$ are randomly initialized, the start model $m=L_{w}(z)$ won’t be $m_{ref}$ without pre-training. In the first stage of DIP-Inv, we update the weights $w$ with the objective function $||m-m_{ref}||^{1}$ to make the start model for the second stage similar to the start model of the conventional method. In the second stage, we load the pre-trained weights first and then update the weights with the objective function 4. Note that the implicit regularization effect does not come from the pre-trained since the reference model is a uniform half-space. Stage one is mainly to save time: The pre-trained weights in the first stage can be stored and reused for future experiments with the same reference model.
@@ II-C Architecture of DIP-Inv
Figure 2: Architecture of DIP-Inv. The structure of hidden blocks 2-4 is shown on the right (hb denotes hidden block). More details are shown in the Appendix.
The structure of the function $L$ is illustrated in Fig 2. The input $z$ is a fixed 1D vector sampled from the Gaussian distribution with mean 0 and standard deviation 10 and is not trainable. The first hidden block is a multi-layer perceptron (MLP) whose output is $y_{1}\in R^{h}$, and the value of $h$ should be adjusted to adapt different mesh designs (i.e. dimension of $M$). Sequentially, $y_{1}$ is reshaped and then fed into 3 hidden blocks where each block has an upsampling layer, a 2D convolutional layer, and a Leaky ReLU activation layer. The last hidden layer only has a 2D convolutional layer followed by a sigmoid activation function. The output of the CNN is cropped to have the same dimension as $M$. Sigmoid activation will squeeze the output into (0, 1). Since we perform the forward modelling on a log-conductivity scale, all elements in the output of the last hidden layer are multiplied by -8 before being fed into the forward simulation. Namely, we assume that the subsurface space has conductivity with range (exp(-8), exp(0)) S/m $\approx$ (3.35e-4, 1) S/m, which covers a large range of natural subsurface materials. This multiplier can be changed to different values to adapt to different conductivity ranges. The dropout layer can also be included in the second stage and we will examine the impact of the regularization effect from the Dropout layer in section IV-C. Although some advanced machine learning models for 2D data use components such as inception blocks, skip connections and residual blocks, they are often based on a CNN architecture. Our DIP-Inv method is also CNN-based and provides a good test for exploring the utility of the implicit regularization effect. As a result, testing this basic model structure is a good first choice in exploring the utility of the implicit regularization effect.

### v2

@@ Abstract
Regularization is critical for solving ill-posed geophysical inverse problems. Explicit regularization is often used, but there are opportunities to explore the implicit regularization effects that are inherent in a Neural Network structure. Researchers have discovered that the Convolutional Neural Network (CNN) architecture inherently enforces a regularization that is advantageous for addressing diverse inverse problems in computer vision, including de-noising and in-painting. In this study, we examine the applicability of this implicit regularization to geophysical inversions. The CNN maps an arbitrary vector to the model space. The predicted subsurface model is then fed into a forward numerical simulation to generate corresponding predicted measurements. Subsequently, the objective function value is computed by comparing these predicted measurements with the observed measurements. The backpropagation algorithm is employed to update the trainable parameters of the CNN during the inversion. Note that the CNN in our proposed method does not require training before the inversion, rather, the CNN weights are estimated in the inversion process, hence this is a test-time learning (TTL) approach. In this study, we choose to focus on the Direct Current (DC) resistivity inverse problem, which is representative of typical Tikhonov-style geophysical inversions (e.g. gravity, electromagnetic, etc.), to test our hypothesis. The experimental results demonstrate that the implicit regularization can be useful in some DC resistivity inversions. We also provide a discussion of the potential sources of this implicit regularization introduced from the CNN architecture and discuss some practical guides for applying the proposed method to other geophysical methods.
@@
Convolutional neural network (CNN), deep image prior (DIP), deep learning (DL), direct-current resistivity (DCR) inversion, and parameterized inversion (PI).
@@ I Introduction
Geophysical inversions are widely used in mining, environmental, and engineering applications [1], [2]. Although forward modelling and optimization strategies have been developed for several decades, it’s still challenging to recover the subsurface structures mainly because this class of problems is highly ill-posed [3]. For this reason, researchers put a lot of effort into exploring well-functioning regularizers for the inversion.
@@
A conventional approach employs the Tikhonov-regularized method where an explicit regularization term, comprised of smallness and smoothness terms, is included in the objective function [1], [3], [4]. The $l_{2}$ norm is standard in the Tikhonov-regularized method [5], but there are works on using different norms to promote sparse or compact structures [4], [6]. Researchers have also developed strategies for incorporating physical properties such as geological or petrophysical data into the Tikhonov-style geophysical geophysical inversions [4], [7], [8]. The recent emergence of deep learning has garnered significant attention from researchers, particularly regarding the integration of machine learning algorithms into the regularization [2], [9], [10], [11].
@@
Supervised learning methods have been employed in scenarios where researchers have prior knowledge of the subsurface geological or petrophysical models. The synthetic datasets can be built based on the prior knowledge to train ML models [12], [13]. One promising approach is to search for a solution in a latent space whose distribution contains prior information learned from the training samples [14]. For example, in Ground-penetrating radar (GPR) inversion, Laloy et al. [15] and Lopez-Alvis et al. [16] trained Variational Autoencoders to map the distributions of the spatial patterns of the subsurface materials into simple distributions and integrated this learned prior into the objective function. In seismic full-waveform inversion (FWI), Mosser et al. [17] trained a generative adversarial network with the knowledge of subsurface geology. These methods predict more realistic subsurface models. The utilization of generative models requires a training dataset, which implies that we need prior knowledge of what we are looking for in the subsurface (e.g. the possible geological models or subsurface materials of the survey domain). Therefore, these approaches utilize a physically motivated prior.
@@
Although the ML models can learn prior knowledge with a synthesis dataset, these methods are not applicable when we have limited prior information. In the past five years, self-supervised and unsupervised learning approaches have been applied to inverse problems. One main approach employs so-called test-time learning (TTL) methods which do not require any training data. Rather than adjusting the weights of a neural network using training samples, the weights of a neural network are learned at the time the neural network is applied to the inversion process. TTL was first applied in the inverse problems in computer vision (CV) [20]. One example of the inverse problem in CV is in-painting (e.g. filling in corrupted portions of an image). This is achieved by learning the self-similarity between the corrupted image and the predicted image. Researchers have found that some neural networks, such as Convolutional Neural Networks (CNNs) and Graph Convolutional Networks (GCNs), inherently enforce a regularization that is advantageous for addressing diverse CV inverse problems or mesh restoration problems [19], [20]. TTL methods have been applied to many biomedical imaging problems[21], [22].
In geophysics, TTL was first adapted to seismic FWI inversion by integrating the ML architecture into the inverse process[23], [24], [25], [26]. He et al. [23] and Wu et al. [24] reparameterized the velocity model into the CNN domain and attributed the regularization effect to the prior information learning by fitting the initial velocity models in the pre-training stage. In contrast, Zhu et al. [26] stated that this regularization effect comes from the CNN architectures as what Ulyanov et al. [20] proposed. Zhu et al. [26] examined this statement in the synthetic cases and found that the implicit regularization effect is robust to noisy data. Instead of having a fixed latent vector as the input in the previous works, researchers have tried replacing the latent vector with the observed seismic data and achieved better inversion results than the conventional FWI with total variation regularization [27]. TTL has also been adapted to joint inversion of geophysical data [28], where researchers used a Bayesian neural network and a CNN to reparametrize the global model parameters and other geological model parameters, respectively. Utilizing trainable weights within the NN to parameterize the subsurface model worked well to regularize the seismic inverse problems, even though the prior they imposed on the solution is not a physically motivated prior.
The physics that governs seismic FWI is the wave equation, and the strategies taken to solve these inverse problems often differ from those used in EM and potential field inversions, which are diffusive in nature. Problems such as EM and potential field inversions are typically solved using a Tikhonov-regularized method where a first-order smoothness term is always included [29]. In this study, we would like to examine the differences between the conventional inversion that uses a smoothness term and an approach that leverages implicit regularization effects from a CNN. We choose the direct current (DC) resistivity inversion, which is one type of EM inverse problem, but the proposed method can be applied to other Tikhonov-style geophysical inversions. To the best of our knowledge, this is the first study that directly investigates the prior captured by the CNN for a DC resistivity inversion by integrating the CNN into the DC resistivity inversion pipeline in a truly unsupervised manner. There are works on applying CNNs in EM inversions [30], [31], [32] in a purely data-driven way. Alternatively, B. Liu et al. [18] proposed the method PhResNet, which reconstructs the subsurface model by a U-net; they trained the U-net with a dataset that has no labels in it. B. Liu et al. [18] stated that using a CNN to parameterize the inverse problem can be beneficial. However, there are two priors in PhResNet: the explicit prior learned from the training dataset and the implicit prior inherently from the CNN structure. It is unclear how the contributions from each of these two priors impact the results since both are intertwined when CNNs are trained with a dataset. Additionally, the generalizability of CNN models in the previous works depends on the size of the training dataset. In contrast, our proposed method does not rely on any training data, so the foundation for our work and previous works are fundamentally different.
@@
We refer to our proposed method as the Deep Image Prior Inversion (DIP-Inv), as it was inspired by the work of [20], which introduced the concept of the Deep Image Prior in computer vision. We will begin with an overview of the methodology and describe the modifications we make to the pipeline of inversion. Then we will show the results of the DIP-Inv. The results illustrate that the CNN reparameterization provides useful regularization effects; therefore, DIP-Inv gives better inversion results than the conventional methods in some cases. After that, we explore different components of the CNN architecture for DIP-Inv and identify the bi-linear upsampling operation, which smooths the final predicted model efficiently in some scenarios, as one of the main sources of this implicit regularization effect.
@@
    Comparing inversion results obtained from conventional Tikhonov-regularized methods with sparse norms and those obtained from the DIP-Inv method on multiple synthetic examples. Our results illustrate that the implicit regularization effect inherently from CNN is beneficial in the Tikhonov-style geophysical inversions.
@@
    Examining some key choices in the CNN architecture to illustrate that the bi-linear upsampling operation is one of the main sources of this implicit regularization effect. And the dropout technique is also useful in some cases.
@@
    Discussing some practical details for connecting the forward simulation, written in SimPEG [33], with PyTorch [34] to parameterize the model with a CNN.
@@ II Methodology
In this section, we will first compare the conventional method and the DIP-Inv method (DIP-Inv) for solving the geophysical inverse problem. Then, we will illustrate the structure of DIP-Inv. Finally, we will explain the objective function of the DIP-Inv method in detail. Note that we only conduct 2D DC resistivity inversions in this study, but this methodology can be adapted to 3D DC resistivity inversions or other Tikhonov-style geophysical inversions.
@@ II-A Overview of the Conventional Inversion Method
$\boldsymbol{\nabla}\cdot\sigma\boldsymbol{\nabla}\phi=-\boldsymbol{\nabla}\cdot\mathbf{j}_{source}$ (1)
where $\sigma$ is the electrical conductivity (which is the reciprocal of electrical resistivity $\rho=1/\sigma$), $\phi$ is the electric potential, and $\mathbf{j}_{source}$ is the source current density. We solve this equation using a finite volume method where we discretize the subsurface using a 2D tensor mesh (c.f. [35]). The mesh consists of a core region, with uniform cells, and padding regions on the lateral boundaries and the bottom of the mesh with expanding cells to ensure boundary conditions are satisfied. We use SimPEG, an open-source package, to implement the forward modelling [33].
@@
The regularization term $\phi_{m}$ generally includes a smallness term, which penalizes the difference from a reference model $m_{ref}$, and a first-order smoothness term that penalizes spatial gradients in the model. The relative influence of each of these terms is weighted by the scalar parameters $\alpha_{s,x,z}$. The standard choice of norms in the regularization is $p_{s,x,z}=2$, however, different norms where $p_{s,x,z}\leq 2$ can be chosen, and we will use the implementation described in [6] for these inversions. Optionally, sensitivity weighting can be included in the weight matrices $W_{s,x,z}$, and these have been shown to help improve the convergence of inversions where $p_{s,x,z}\leq 2$, [6].
@@
The inversion aims to recover a model that acceptably fits the data. With the assumption that noise is Gaussian and appropriate uncertainties have been used to define $W_{d}$ [29], an acceptable fit is often taken to be $\phi_{d}\approx\phi_{d}^{*}$ where $\phi_{d}^{*}=N/2$. For all of the conventional inversions that we consider in this paper, a $\beta$-cooling strategy will be adopted, which is a common practice [29]. We first estimate an initial $\beta$ by comparing the magnitude of $\phi_{d}$ and $\phi_{m}$ when evaluated on a random model and set $\beta_{0}=\beta_{s}\phi_{d}/\phi_{m}$ to ensure that the regularization has the most influence early in the inversion.
We then employ an optimization method, commonly a second-order approach such as inexact Gauss Newton, to compute an update to the model [5]. We then reduce $\beta$, reducing the influence of the regularization and allowing the data misfit term to have more influence in the solution and again compute an update to the model. In an inversion where $l_{2}$ norms are used in the regularization, this process is stopped when we have reached the target misfit. If using a different choice of norms in the smallness or smoothness terms of the regularization (e.g. $p_{s},p_{x},p_{z}\leq 2$), then we follow the methodology described in [6]. First, we solve the inverse problem using $p_{s},p_{x},p_{z}=2$ and then use this model as a starting model for the optimization problem with $p_{s},p_{x},p_{z}\leq 2$. Since the regularization is no longer convex when norms are less than 2, this is a more computationally challenging problem. To address this, the $\ell_{p}$ norm is approximated using a Lawson Norm [36] that is evaluated and updated using the Iterative Reweighted Least-Squares (IRLS) algorithm [6]. Additionally, sensitivity weighting can optionally be included [37], [38].
[…]
Figure 1: (a) is the true conductivity model. (b) is the result of a $l_{2}$ inversion using only smallness. (c) is the result of a $l_{2}$ inversion using smallness and smoothness. (d) is the result of a $l_{1}$ inversion using smallness and smoothness. (e) is the result of a $l_{1}$ inversion using smallness, smoothness, and sensitivity weighting.
II-B Overview of the Proposed Parameterized Inversion Method: DIP-Inv
Figure 2: Proposed inversion pipeline. DIP-Inv reparameterizes the mesh space into the CNN-weights space.
We are interested in exploring if the implicit regularization can be advantageous in solving the Tikhonov-style geophysical inversions, so the main modification that we make to the inversion pipeline posed in 2 is that we parameterize the model in terms of a CNN-based function.
In our DIP-Inv method, a CNN-based function $L$, whose weights are denoted as $w$, maps an arbitrary vector $z\in R^{k}$ to a model $m$ in the mesh space $M\in R^{W\times H}$ [Fig. 2].
@@ II-B Overview of the Proposed Parameterized Inversion Method: DIP-Inv
Note that the padding cells are included in $m$. The predicted subsurface model $m$ is then fed into the forward numerical simulation implemented by SimPEG to generate the corresponding predicted measurements $d^{pred}$. The modified optimization problem that we aim to solve is
@@
Compared with the supervised end-to-end machine learning methods that aim to remove the physical forward simulations in the standard EM inversion, DIP-Inv imposes physical constraints in the forward simulation. Compared with the conventional Tikhonov-regularized methods, the objective function in the DIP-Inv method doesn’t have the smoothness term since the smoothness constraint is implicitly included in the upsampling operator in $L$ as will be discussed in section IV-A. Also, we don’t use the sensitivity weighting. The backpropagation algorithm, which is implemented by PyTorch, is employed to update the trainable parameters $w$ in the function $L$.
Since the weights $w$ are randomly initialized, the starting model $m=L_{w}(z)$ won’t be $m_{ref}$ without pre-training. In the first stage of DIP-Inv, we update the weights $w$ with the objective function $||m-m_{ref}||^{1}$ to make the starting model for the second stage similar to the starting model of the conventional method. In the second stage, we load the pre-trained weights and then update the weights with the objective function 4. Note that the implicit regularization effect does not come from the pre-training since the reference model is a uniform half-space. The first stage is mainly for saving time: the pre-trained weights in the first stage can be stored and reused for future experiments with the same reference model.
@@ II-C Architecture of DIP-Inv
The structure of the function $L$ is illustrated in Fig. 3. The input $z$ is a fixed 1D vector sampled from the Gaussian distribution with mean 0 and standard deviation 10 and is not trainable. The first hidden block is a multi-layer perceptron (MLP) whose output is $y_{1}\in R^{h}$, and the value of $h$ should be adjusted to adapt different mesh designs (i.e. dimension of $M$). Subsequently, $y_{1}$ is reshaped and then fed into 3 hidden blocks where each block has an upsampling layer, a 2D convolutional layer, and a Leaky ReLU activation layer. The last hidden layer only has a 2D convolutional layer followed by a sigmoid activation function. The output of the CNN is cropped to have the same dimension as $M$. The Sigmoid activation will squeeze the output into (0, 1).
Figure 3: Architecture of DIP-Inv. The structure of hidden blocks 2-4 is shown on the right (hb denotes hidden block). More details are shown in Table. I.
TABLE I: Details of the CNN architecture implemented in vanilla DIP-Inv
hb1 Fully-connected layer (8, 174)
LeakyReLU (negative slope 0.2)
reshape (1,1,6,29)
hb2 Bi-linear upsampling ($k=2$)
Convolutional ((1,64), kernel ($3\times 3$), stride (1), padding (0))
LeakyReLU (negative slope 0.2)
hb3 Bi-linear upsampling ($k=2$)
Convolutional ((64, 32), kernel ($3\times 3$), stride (1), padding (0))
LeakyReLU (negative slope 0.2)
hb4 Bi-linear upsampling ($k=2$)
Convolutional ((32, 8), kernel ($3\times 3$), stride (1), padding (0))
LeakyReLU (negative slope 0.2)
hb5 Convolutional ((8, 1), kernel ($3\times 3$), stride (1), padding (0))
Sigmoid
output cropped, $\times$ scale factor (problem-specific, -8 in this study)
The output of the last hidden layer is on a log-conductivity scale. All elements in the output are multiplied by -8 before being fed into the forward simulation, so the subsurface space can have conductivity values with range (exp(-8), exp(0)) S/m $\approx$ (3e-4, 1) S/m, which covers a large range of natural subsurface materials. This multiplier can be changed to different values to adapt to different conductivity ranges.
The dropout layer can also be included in the second stage and we will examine the impact of the regularization effect from the dropout layer in section IV-C. Although more advanced models have components such as inception blocks, skip connections or residual blocks, they are often based on a CNN architecture. As a result, this basic structure of CNN employed by DIP-Inv provides a good test for exploring the utility of the implicit regularization effect.